Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
lumaca, nardecchia, sposito, taranto - 11:31 Saturday 03 August 2024 (64851) Print this report
Reversal of WI outer ring horizontal movement

This morning, due to the issues in the lock acquisitions (64847), the WI outer ring movement performed yesterday (64844) has been reverted. To do that, Gianmatteo switched ON the WI picomotors in the TCS room. 
The movement (BS_REC TY 1.5 cm backward (5556+2778 steps)) was completed at 09.23 UTC with the arms locked on IR. 

Comments to this report:
mwas - 8:13 Monday 05 August 2024 (64865) Print this report

The change in WI CO2 alignment has clearly improved the situation. The mean value of B1s during the steps when the OMC is locked has been reduced from ~40mW to ~25mW. Keeping well below the 45mW threshold that triggers the OMC safety and prevents the lock acquisition.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×